Present age ratio of father and his son is 2:1.Before 10 years age ratio of them is 5:2.
What is the present age of son?

Answer by Cromlix(4381)

hi there,
Present age ratio of father and his son is 2:1.
F/S = 2/1
Cross multiply
F = 2S
Before 10 years age ratio of them is 5:2.
F - 10/S - 10 = 5/2
Cross multiply
5S - 50 = 2F - 20
Replace F with 2S from first equation
5S - 50 = 2(2S) - 20
Multiply out
5S - 50 = 4S - 20
Collect like terms
5S - 4S = - 20 + 50
.........S = 30
Present age of son = 30 years old
Hope this helps:-)